itzmeee:

2x-y=8 x+2y=4 For the system of equations above, what is the value of x+y.

YourlocalRandom2000:

x + y = 4 + 0 x + y = 4

maccinkylie:

first one is 2 x 5 then y = 2

maccinkylie:

x=5 y+2